OpenStudy (anonymous):

help me get the general solution please. y'' - 3y' -4y=30e^{4x}

OpenStudy (anonymous):

D2-3D-4=0 -> D=-1 , 4 so e-x & e4x are homo solutions. Yp=Axe4x then find a

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A=10 so y=c1e-x+c2e4x+10xe4x

OpenStudy (anonymous):

no c1-5=c !!!

OpenStudy (anonymous):

looks like college math to me

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

what class is that

OpenStudy (anonymous):

diffeerential equation

OpenStudy (anonymous):

took is 2 yrs ago but i will try to help you, where are you stuck?

OpenStudy (anonymous):

in determining the coefficient . i keep on getting zero.